jigglypuff314 (jigglypuff314):

how to convert ? 6L (nitrogen) at 2500psig would equal how many L at 0psig?

jigglypuff314 (jigglypuff314):

I have no idea what any of this means btw :P my dad needs to know for work or something...

jigglypuff314 (jigglypuff314):

@wio could you help pwease? :3

OpenStudy (31356):

Sorry, I have no idea about this, wish I could help. :D

OpenStudy (31356):

What's psig?

jigglypuff314 (jigglypuff314):

pounds per square inch... 0.o

OpenStudy (31356):

Nice music on Soundcloud too :D

jigglypuff314 (jigglypuff314):

thanks :)

OpenStudy (31356):

Oh okay, my pleasure, what is your dad trying to measure?

jigglypuff314 (jigglypuff314):

idk I think he has compressed nitrogen in a tank and he want to know the volume of it at normal pressure...

OpenStudy (31356):

Oh okay, good luck with your dad on that.

jigglypuff314 (jigglypuff314):

lol thanks ^_^

OpenStudy (31356):

:D Any time

jigglypuff314 (jigglypuff314):

can anyone help me pwaese? :3 @whpalmer4 @jdoe0001 @ganeshie8

OpenStudy (jdoe0001):

hmm dunno :(

jigglypuff314 (jigglypuff314):

@Mertsj if you're not too busy, could you help? or if you don't know either, could you tag someone that might know how to help me?

OpenStudy (mertsj):

I think you want the gas law: \[P _{1}V _{1}=P _{2}V _{2}\]

jigglypuff314 (jigglypuff314):

ok, thanks! :)

jigglypuff314 (jigglypuff314):

would it be 1011.708 L then?

jigglypuff314 (jigglypuff314):

idk cuz normal pressure for psig is 0 right?

OpenStudy (mertsj):

Here comes the cavalry

OpenStudy (anonymous):

Since you can't divide by 0, it's likely you'll need to convert psig.

OpenStudy (whpalmer4):

at sea level, the atmospheric pressure is 14.7 psi. he's got a psig reading, which doesn't account for the sea level pressure — so his actual pressure is 2500 psi + 14.7 psi. \[P_1 = 2514.7\]\[V_1=6\]\[P_2=14.7\]\[P_1V_1 = P_2V_2\]\[V_2=\frac{P_1V_1}{P_2} = \frac{2514.7*6}{14.7} = 1026.4 \]

jigglypuff314 (jigglypuff314):

ohhh okay, got it now! :) thanks! :D

OpenStudy (whpalmer4):

This answer comes with no warranty of suitability for filling breathing tanks, hot air balloons, high pressure boilers, etc. :-)

jigglypuff314 (jigglypuff314):

lol my dad's reaction: "ok, so 1026.4L in 46 hours would be about 22L then two machines using it... so about 11L per machine per hour... 0.o that can be right... ... *** there must be a leak then... "